# DESI constraints on $\alpha$-attractor inflationary models

### v2

@@ Abstract
The recent results on the baryon acoustic oscillations measurements from the DESI collaboration have shown tantalizing hints for a time-evolving dark energy equation of state parameter $w(z)$, with a statistically significant deviation from the cosmological constant and cold dark matter ($\Lambda$CDM) model. One of the simplest and theoretically well-motivated plausible candidates to explain the observed behavior of $w(z)$, is scalar-field quintessence. Here, we consider a class of models known as $\alpha$-attractor, which describe in a single framework both inflation and the late-time acceleration of the Universe. Using the recent DESI data, in conjunction with other cosmological observations, we place stringent constraints on $\alpha$-attractor models and compare them to the $\Lambda$CDM  model. We find the $\alpha$ parameter of the theory, which is physically motivated from supergravity and supersymmetry theories to have the values $3\alpha\in\{1,2,3,4,5,6,7\}$, is constrained to be $\alpha\simeq 1.89_{-0.35}^{+0.40}$. In addition, we find that the rest of the cosmological parameters of the model agree with the corresponding values of $\Lambda$CDM, while a bayesian analysis finds strong support in favor of the $\alpha$-attractor model. We also highlight an interesting connection between the $\alpha$-attractor models and the stochastic gravitational wave background, where a contribution to the latter could derive from an enhancement of inflationary gravitational waves at high frequencies due to an early kination phase, thus providing an interesting alternative way to constrain the theory in the near future.
@@ I Introduction
Still, despite the success of this simple model, several tensions have recently appeared between low redshift and high redshift probes, see for example Ref. Perivolaropoulos and Skara (2022). Moreover, the cosmological constant also causes fine-tuning as its value differs by orders of magnitude with what is predicted by quantum field theories Weinberg (1989); Carroll (2001), faces various theoretical problems Martin (2012), and is also plagued by the so-called coincidence problem Zlatev et al. (1999). These issues stimulated the creation of several DE models based on ad-hoc modification of gravity or scalar fields that mediate the force between particles, e.g. canonical scalar fields Ratra and Peebles (1988); Wetterich (1988); Caldwell et al. (1998), scalar fields with generalized kinetic terms Armendariz-Picon et al. (2000, 2001); Chiba et al. (2000), non-minimal couplings Uzan (1999); Perrotta et al. (1999); Riazuelo and Uzan (2002); Chiba (1999) or coupled DE models Dent et al. (2009), in addition to GR.
@@
In the most general picture of single-field slow-roll inflation, the inflaton evolves along a flat region of its potential, slowly rolling and driving cosmic acceleration. As the inflaton exits this flat region, it begins to move faster, leading to the end of inflation. While single-field slow-roll inflation successfully addresses classical cosmological problems, several intriguing features emerge when moving beyond this simplified model. These include distinguishing effective field theories within the string landscape Achúcarro and Palma (2019), producing primordial black holes without having a strong scale dependence at small scales Palma et al. (2020), and possible violations of the consistency relations for local primordial non-Gaussianity Cabass et al. (2022).
@@
Cosmological observations are particularly well-suited to constrain these models, as both early- and late-time physics can leave imprints in the LSS and the CMB (both measured by dedicated surveys), thus allowing us to probe for hints of these models. In this regard, the Dark Energy Spectroscopic Instrument (DESI) is particularly useful for probing the physics of DE at late times, by mapping the LSS of the Universe. Recently, the collaboration released the first year data Adame et al. (2024a, b, c), where they report the latest results from BAO measurements and the resulting cosmological constraints. Interestingly, a statistically significant deviation from $\Lambda$CDM was also reported, which could potentially be explained by a quintessence model Tada and Terada (2024); Yin (2024); Bhattacharya et al. (2024); Ramadan et al. (2024); Wolf et al. (2024a); Giarè et al. (2024a).
In this paper, we use the latest observational CMB and LSS data, which provide information on early- and late-time physics respectively, to place constraints on the $\alpha$-attractor models and compare them to the $\Lambda$CDM, model. This allows us to test the compatibility of these models with the latest cosmological data and provide a unified explanation for both inflation and DE. In this context, we also explore the connections between the $\alpha$-attractor model and gravitational waves (GWs) Wali Hossain et al. (2015); Dimopoulos and Owen (2017); Ahmad et al. (2019); Wang et al. (2024). A key feature of quintessential inflation is the post-inflation phase dominated by the scalar field’s kinetic energy. This alters the Hubble evolution, enhancing inflationary GWs at high frequencies compared to the standard radiation-dominated Universe Giovannini (1998); Peebles and Vilenkin (1999); Giovannini (1999a, b, 2009); Tashiro et al. (2004); Figueroa and Tanin (2019a, b); Duval et al. (2024). To investigate this, we first estimate the amplitude of the GW spectrum based on the parameter values inferred from CMB and LSS data. We then discuss potential constraints on the parameter space, including limits from big-bang nucleosynthesis (BBN) on the GW amplitude at high frequencies, as well as theoretical considerations regarding the allowed duration of the kination phase, which is related to the reheating temperature.
@@ II $\alpha$-attractor models
More conveniently, one can redefine the scalar field in terms of a canonically normalized field $\varphi$, as
@@
When $\alpha\rightarrow\infty$ then the two fields, namely $\phi$ and $\varphi$ become equal, while in general in terms of $\varphi$, the potential becomes sufficiently flat to support accelerated expansion. It is possible to consider different functions to build up the inflaton potential, but it is interesting to focus on functions that allow for the presence of an inflection point.
Following the discussion in Ref. Akrami et al. (2021, 2018); Dimopoulos and Owen (2017), we can select the following potentials of exponential form,
$V(\phi)$ $=$ $M^{2}e^{\gamma\left(\frac{\phi}{\sqrt{6\alpha}}-1\right)}+V_{0},$ (3)
@@
In this analysis, we set $V_{0}=0$. In the context of $\alpha$-attractor models, $\alpha$ is a free parameter that can be constrained phenomenologically. Nevertheless, interesting theoretical values of $3\,\alpha\in\{1,2,3,4,5,6,7\}$, are set by supersymmetry and supergravity, as reported in Refs. Ferrara and Kallosh (2016); Kallosh et al. (2017a, b).
On the tails of the potentials for large and positive $\varphi$, where we assume inflation to take place, the form of the effective potential is approximated by the expression
@@
With this approximated form, the number of e-folds $N_{\ast}$ is approximately given by Akrami et al. (2021)
$N_{\ast}\simeq\frac{3\alpha}{4\gamma}e^{-{\frac{2\varphi_{\ast}}{\sqrt{6\alpha}}}}\,,$ (6)
where $\varphi_{\ast}$ is the value of the scalar field when the corresponding mode exits the Hubble radius. Using this, $M$ is corresponded to $\mathcal{A}_{s}$, the amplitude of the primordial scalar power spectrum, as
@@
where $\rho_{\mathrm{m}}$ and $\rho_{\mathrm{r}}$ are the matter (baryons $\rho_{\mathrm{b}}$ plus cold dark matter $\rho_{\mathrm{cdm}}$) and radiation energy densities, respectively, primes are derivatives with respect to the number of e-folds $N=\ln(a)$, $H=\dot{a}/a$ is the Hubble parameter, and a dot is a derivative with respect to the cosmic time $t$. The scalar field freezes after reheating and begins to move again once it overcomes Hubble friction. To calculate the late-time dynamics of the scalar field, we set the initial condition to $\varphi_{\mathrm{init}}=-10$.
@@
Following Ref. Akrami et al. (2021), we focus on the effects of the $\alpha$-attractor models at late times, assuming the early and late-time physics are connected via the relations given by Eqs. (7)-(8). Then, we modify the publicly available Boltzmann code CLASS Lesgourgues (2011); Blas et al. (2011), which already includes a module for solving the quintessence equations at late times, and we include the potential of Eq. (4) along with the parameters that connect early and late times via Eqs. (7)-(8). For our analysis, we set $V_{0}=0$ in Eq. (4). It should be noted that the parameter $\gamma$ that appears in Eq. (4) is not free to vary, but instead is determined and automatically adjusted using shooting, such that the Friedmann equation Eq. (11) is satisfied today, i.e. $H(a=1)=H_{0}$.
@@ III Observational data and methodology
In our analysis, we use three datasets comprising of CMB, BAO, and supernovae measurements. More specifically, we use the full Planck 2018 data, i.e. the TT,TE,EE,lowE and Lensing likelihoods Aghanim et al. (2020), hereby referred to as CMB$+$Lensing, and the newly released DESI DR1 BAO measurements Adame et al. (2024b) (for an alternative study on BAO measurements see Anselmi et al. (2023); O’Dwyer et al. (2020); Anselmi et al. (2019)). Lastly, we also use the Pantheon$+$ Type Ia supernovae (SnIa) data Riess et al. (2022).
We consider two distinct cases: the vanilla flat $\Lambda$CDM  model, which is characterized by five key cosmological parameters $\{\omega_{b},\Omega_{\mathrm{m}},n_{s},\mathcal{A}_{s},H_{0}\}$ with $\omega_{\mathrm{b}}\equiv\Omega_{\mathrm{b}}h^{2}$, and the $\alpha$-attractor model, which has six key parameters $\{\omega_{\mathrm{b}},\Omega_{\mathrm{m}},n_{s},\alpha,\mathcal{A}_{s},H_{0}\}$. The uncertainties in other free parameters, such as the optical depth $\tau$, are marginalized over in the results presented in the next section. We consider reasonable priors for all the standard $\Lambda$CDM parameters to encompass a broad parameter space, while we choose $\alpha\in[0.01,5]$ and $\alpha\in[0.5,3.5]$ for the CMB$+$Lensing and the full data compilation respectively.
@@ IV MCMC analysis
Finally, for the results that take into account the full data combination, we compute the Bayesian evidences (marginal likelihoods) of the $\alpha$-attractor and $\Lambda$CDM  models using the MCEvidence package Heavens et al. (2017), see Table 3. This algorithm obtains the posterior for the marginal likelihood, using the $k$-th nearest-neighbour Mahalanobis distance Mahalanobis (1936) in the parameter space. In this analysis, we consider the $k=1$ case to minimize the effects of the inaccuracies associated with larger dimensions of the parameter space and smaller sample sizes.
Models $\ln(B)$ $\Delta\ln(B)$ $\Delta\chi^{2}$
@@
Table 3: The log Bayes evidences and the differences in $\chi^{2}$ for the two models, using the combination of CMB, Lensing, DESI and Pantheon$+$ data.
To evaluate the strength of evidence for or against a model in a comparison between two models using the Bayes factor $B$ (the ratio of evidences), the revised Jeffreys’ scale can be applied Trotta (2008). According to this scale, if $|\ln B|<1$, the models are similar, with neither being particularly favored. When $2.51<|\ln B|<$ 2.5, there is weak evidence supporting one model. For $52.5<|\ln B|<$ 5, the evidence becomes moderate, and if $|\ln B|>5$, there is strong evidence favoring one model over the other.
As can be seen in Table 3, we find $\Delta\ln(B)\sim 6.36$ which implies strong evidence in favor of the $\alpha$-attractor model against $\Lambda$CDM. We note that this is in agreement with the results found by DESI Adame et al. (2024b), namely that $\Lambda$CDM  is not preferred by the CMB+Pantheon+DESI combination. While this deviation from a constant equation of state could be attributed to new physics Seto and Toda (2024); Chudaykin and Kunz (2024); Ye et al. (2024); Wang and Piao (2024); Wolf et al. (2024b); Li et al. (2024), it might also be due to systematics in the data Liu et al. (2024); Pérez-Fernández et al. (2024); Krolewski et al. (2024).

### v3

@@ Abstract
The recent results on the baryon acoustic oscillations measurements from the DESI collaboration have shown tantalizing hints for a time-evolving dark energy equation of state parameter $w(z)$, with a statistically significant deviation from the cosmological constant and cold dark matter ($\Lambda$CDM) model. One of the simplest and theoretically well-motivated plausible candidates to explain the observed behavior of $w(z)$, is scalar-field quintessence. Here, we consider a class of models known as $\alpha$-attractor, which describe in a single framework both inflation and the late-time acceleration of the Universe. Using the recent DESI data, in conjunction with other cosmological observations, we place stringent constraints on $\alpha$-attractor models and compare them to the $\Lambda$CDM  model. We find the $\alpha$ parameter of the theory, which is physically motivated from supergravity and supersymmetry theories to have the values $3\alpha\in\{1,2,3,4,5,6,7\}$, is constrained to be $\alpha\simeq 1.89_{-0.35}^{+0.40}$. In addition, we find that the rest of the cosmological parameters of the model agree with the corresponding values of $\Lambda$CDM, while a Bayesian analysis finds strong support in favor of the $\alpha$-attractor model. We also highlight an interesting connection between the $\alpha$-attractor models and the stochastic gravitational wave background, where a contribution to the latter could derive from an enhancement of inflationary gravitational waves at high frequencies due to an early kination phase, thus providing an interesting alternative way to constrain the theory.
@@ I Introduction
Still, despite the success of this simple model, several tensions have recently appeared between low redshift and high redshift probes, see for example Ref. Perivolaropoulos and Skara (2022). Moreover, the cosmological constant also suggests fine-tuning as its value differs by orders of magnitude with what is predicted by quantum field theories Weinberg (1989); Carroll (2001), faces various theoretical problems Martin (2012), and is also plagued by the so-called coincidence problem Zlatev et al. (1999). Specifically, Ref. Martin (2012) discusses several aspects of the cosmological constant model, not only in the original version discussed in Ref. Weinberg (1989), and that in reality the renormalized value of the zero-point energy density is in fact far from being 122 orders of magnitude larger than the critical energy density, as often mentioned in the literature. These issues stimulated the creation of several DE models based on ad-hoc modification of gravity or scalar fields that mediate the force between particles, e.g. canonical scalar fields Ratra and Peebles (1988); Wetterich (1988); Caldwell et al. (1998), scalar fields with generalized kinetic terms Armendariz-Picon et al. (2000, 2001); Chiba et al. (2000), non-minimal couplings Uzan (1999); Perrotta et al. (1999); Riazuelo and Uzan (2002); Chiba (1999) or coupled DE models Dent et al. (2009), in addition to GR.
@@
In the most general picture of single-field slow-roll inflation, the inflaton evolves along a flat region of its potential, slowly rolling and driving cosmic acceleration. As the inflaton exits this flat region, it begins to move faster, leading to the end of inflation. While single-field slow-roll inflation successfully addresses classical cosmological problems, several intriguing features emerge when moving beyond this simplified model. These include distinguishing effective field theories within the string landscape Achúcarro and Palma (2019), producing primordial black holes without having a strong scale dependence at small scales Palma et al. (2020); Fumagalli et al. (2023), and possible violations of the consistency relations for local primordial non-Gaussianity Cabass et al. (2022).
@@
Cosmological observations are particularly well-suited to constrain these models, as both early- and late-time physics can leave imprints in the LSS and the CMB (both measured by dedicated surveys), thus allowing us to probe for hints of these models. In this regard, the Dark Energy Spectroscopic Instrument (DESI) is particularly useful for probing the physics of DE at late times, by mapping the LSS of the Universe. Recently, the collaboration released the first year data Adame et al. (2024a, b, c), where they report the latest results from BAO measurements and the resulting cosmological constraints. Interestingly, a statistically significant deviation from $\Lambda$CDM was also reported, which could potentially be explained by a quintessence model Tada and Terada (2024); Yin (2024); Bhattacharya et al. (2024); Ramadan et al. (2024); Wolf et al. (2024a); Giarè et al. (2024a) or systematics in the DESY3 data Roy Choudhury and Okumura (2024).
In this paper, we use the latest observational CMB and LSS data, which provide information on early- and late-time physics respectively, to place constraints on the $\alpha$-attractor models and compare them to the $\Lambda$CDM, model. This allows us to test the compatibility of these models with the latest cosmological data and provide a unified explanation for both inflation and DE. In this context, we also explore the connections between the $\alpha$-attractor model and gravitational waves (GWs) Wali Hossain et al. (2015); Dimopoulos and Owen (2017); Ahmad et al. (2019); Wang et al. (2024). A key feature of quintessential inflation is the post-inflation phase dominated by the scalar field’s kinetic energy. This alters the Hubble evolution, enhancing inflationary GWs at high frequencies compared to the standard radiation-dominated Universe Giovannini (1998); Peebles and Vilenkin (1999); Giovannini (1999a, b, 2009); Tashiro et al. (2004); Figueroa and Tanin (2019a, b); Duval et al. (2024). To investigate this, we first estimate the amplitude of the GW spectrum based on the parameter values inferred from CMB and LSS data. We then discuss constraints on these parameters using the current upper bounds on the GW amplitude from big-bang nucleosynthesis (BBN) and the CMB, as well as theoretical limits related to the allowed duration of the kination phase, which is tied to the reheating temperature.
@@ II $\alpha$-attractor models
A key feature of $\alpha$-attractor models is that the kinetic term in the original Lagrangian has a pole at $\phi=\pm\sqrt{6\alpha}$. One can redefine the scalar field in terms of a canonically normalized field $\varphi$, as
@@
With this transformation, this pole is shifted to infinity at the cost of introducing an inflection point in the potential. When $\alpha\rightarrow\infty$ then the two fields, namely $\phi$ and $\varphi$ become equal, while in general in terms of $\varphi$, the potential becomes sufficiently flat to support accelerated expansion. It is possible to consider different functions to build up the inflaton potential, but it is interesting to focus on functions that allow for the presence of an inflection point.
Based on the discussion in Ref. Akrami et al. (2021, 2018); Dimopoulos and Owen (2017), we consider the following exponential potential, which naturally arises in high-energy theories such as supergravity or string theory:
$V(\phi)=M^{2}e^{\gamma\left(\frac{\phi}{\sqrt{6\alpha}}-1\right)}+V_{0}.$ (3)
By applying the transformation of Eq. (2), we obtain
@@
In our analysis, we set $V_{0}=0$. In the context of $\alpha$-attractor models, $\alpha$ is a free parameter that can be constrained phenomenologically. Nevertheless, interesting theoretical values of $3\,\alpha\in\{1,2,3,4,5,6,7\}$, are set by supersymmetry and supergravity, as reported in Refs. Ferrara and Kallosh (2016); Kallosh et al. (2017a, b).
In the asymptotic regime, for large and positive $\varphi$, where we assume inflation occurs, the effective potential is approximated by
@@
[…]
$N_{\ast}\simeq\frac{3\alpha}{4\gamma}e^{-{\frac{2\varphi_{\ast}}{\sqrt{6\alpha}}}}\,.$ (6)
Using this, $M$ can be related to $\mathcal{A}_{s}$, the amplitude of the primordial scalar power spectrum, as
@@
where $\rho_{\mathrm{m}}$ and $\rho_{\mathrm{r}}$ are the matter (baryons $\rho_{\mathrm{b}}$ plus cold dark matter $\rho_{\mathrm{cdm}}$) and radiation energy densities, respectively, primes are derivatives with respect to the number of e-folds $N=\ln(a)$, $H=\dot{a}/a$ is the Hubble parameter, and a dot is a derivative with respect to the cosmic time $t$.
The scalar field freezes after reheating and begins to move again once it overcomes Hubble friction. To calculate the late-time dynamics of the scalar field, we set the initial condition to $\varphi_{\mathrm{init}}=-10$. Note that the value of the initial $\phi$ is not very important, one just needs to make sure the field starts rolling after the freeze epoch and before the DE domination phase. For comparison reasons we used the same value as in Ref. Akrami et al. (2021).
@@
Following Ref. Akrami et al. (2021), we focus on the effects of the $\alpha$-attractor models at late times, assuming the early and late-time physics are connected via the relations given by Eqs. (7)-(8). Then, we modify the publicly available Boltzmann code CLASS Lesgourgues (2011); Blas et al. (2011), which already includes a module for solving the quintessence equations at late times, and we include the potential of Eq. (4) along with the parameters that connect early and late times via Eqs. (7)-(8). For our analysis, we set $V_{0}=0$ in Eq. (4).
It should be noted that the parameter $\gamma$ that appears in Eq. (4) is not free to vary, but instead is determined and automatically adjusted using shooting, such that the Friedmann equation Eq. (11) is satisfied today, i.e., $H(a=1)=H_{0}$. The reason for employing the shooting method is that, in an MCMC approach, a point is selected a priori in the parameter space, and the process jumps to that point where we calculate the observables using the CLASS code. During this process, one has to solve both the Friedmann and Klein-Gordon equations simultaneously, and the condition $H(a=1)/H_{0}=1$ might not be satisfied for the given set of parameter values. Consequently, we need to transform the boundary value problem, which has conditions set at early times and requires $H(a=1)/H_{0}=1$, into a pure initial condition problem that inherently satisfies all the required conditions. In our case, we have to adjust one of the parameters of the model, either $\Omega_{\rm m}$ or $\alpha$ or $\gamma$. Given that $\gamma$ is the least physically motivated parameter, we select it for the shooting.
@@ III Observational data and methodology
In our analysis, we use three datasets comprising of CMB, BAO, and supernovae measurements. More specifically, we use the full Planck 2018 data, i.e., the TT,TE,EE,lowE and Lensing likelihoods Aghanim et al. (2020), hereby referred to as CMB$+$Lensing, and the newly released DESI DR1 BAO measurements Adame et al. (2024b) (for an alternative study on BAO measurements see Anselmi et al. (2023); O’Dwyer et al. (2020); Anselmi et al. (2019)). Lastly, we also use the Pantheon$+$ Type Ia supernovae (SnIa) data Riess et al. (2022).
We consider two distinct cases: the vanilla flat $\Lambda$CDM  model, which is characterized by five key cosmological parameters $\{\omega_{b},\Omega_{\mathrm{m}},n_{s},\mathcal{A}_{s},H_{0}\}$ with $\omega_{\mathrm{b}}\equiv\Omega_{\mathrm{b}}h^{2}$, and the $\alpha$-attractor model, which has six key parameters $\{\omega_{\mathrm{b}},\Omega_{\mathrm{m}},n_{s},\alpha,\mathcal{A}_{s},H_{0}\}$. In the $\alpha$-attractor model, the theory is characterized by three fundamental physical parameters: $\alpha$, $M$, and $N_{*}$. Here, we treat $\alpha$ as a free parameter to be constrained, while $M$ and $N_{*}$ can be related to the two CMB parameters, $n_{s}$ and $\mathcal{A}_{s}$, via Eqs. (7) and (8). Although parameter constraints could be expressed in terms of the fundamental physical parameters, for consistency and direct comparison with the $\Lambda$CDM model, we adopt the latter parameter set. As a result, the $\alpha$-attractor model can be viewed as a one-parameter extension of the $\Lambda$CDM framework, with $n_{s}$ and $\mathcal{A}_{s}$ implicitly encoding information about $M$ and $N_{*}$.
The uncertainties in other free parameters, such as the optical depth $\tau$ and the Planck nuisance parameters, which are non-cosmological and related to the instrument, are marginalized over in the results presented in the next section. We consider reasonable priors for all the standard $\Lambda$CDM parameters to encompass a broad parameter space, while we choose $\alpha\in[0.01,5]$ and $\alpha\in[0.5,3.5]$ for the CMB$+$Lensing and the full data compilation respectively.
@@ IV MCMC analysis
Finally, for the results that take into account the full data combination, we compute the Bayesian evidence (marginal likelihoods) of the $\alpha$-attractor and $\Lambda$CDM  models using the MCEvidence package Heavens et al. (2017), see Table 3. This algorithm obtains the posterior for the marginal likelihood, using the $k$-th nearest-neighbor Mahalanobis distance Mahalanobis (1936) in the parameter space. In this analysis, we consider the $k=1$ case to minimize the effects of the inaccuracies associated with larger dimensions of the parameter space and smaller sample sizes.
Models $\ln E_{i}$ $\ln B$ $\Delta\chi^{2}$
@@
Table 3: The log-evidence $\ln E_{i}$, the log Bayes $\ln(B)$, where $B\equiv E_{i}/E_{j}$, and the differences in $\chi^{2}$ for the two models, using the combination of CMB, Lensing, DESI and Pantheon$+$ data.
To evaluate the strength of evidence for or against a model in a comparison between two models using the Bayes factor $B\equiv E_{i}/E_{j}$ (the ratio of evidence between the two models), the revised Jeffreys’ scale can be applied Trotta (2008). According to this scale, if $|\ln B|<1$, the models are similar, with neither being particularly favored. When $2.51<|\ln B|<$ 2.5, there is weak evidence supporting one model. For $52.5<|\ln B|<$ 5, the evidence becomes moderate, and if $|\ln B|>5$, there is strong evidence favoring one model over the other.
As can be seen in Table 3, we find $\ln B\sim 6.36$ which implies strong evidence in favor of the $\alpha$-attractor model against $\Lambda$CDM. We note that this is in agreement with the results found by DESI Adame et al. (2024b), namely that $\Lambda$CDM  is not preferred by the CMB+Pantheon+DESI combination. While this deviation from a constant equation of state could be attributed to new physics Seto and Toda (2024); Chudaykin and Kunz (2024); Ye et al. (2024); Wang and Piao (2024); Wolf et al. (2024b); Li et al. (2024), it might also be due to systematics in the data Liu et al. (2024); Pérez-Fernández et al. (2024); Krolewski et al. (2024).
An interesting point to note is that while the $\Delta\chi^{2}$ between the $\alpha$-attractor model and $\Lambda$CDM is only $\Delta\chi^{2}\sim-0.8$, the log-Bayes is significantly larger at $\ln B\sim 6.36$. To eliminate any possible issues with numerical instabilities, we further employed both a custom code and the MontePython minimizer function to estimate the minimum $\chi^{2}$, confirming that our results are robust. To understand the cause of having similar chi-squares but significantly different log-Bayes values, we may use the Savage-Dickey formula; for a derivation, see Appendix C of Ref. Arjona et al. (2024). In this case, we consider $\ln B$ for the two nested models: a simpler model $M^{\prime}$ with $n^{\prime}$ parameters and a more general model $M$ with $n$ parameters (such that $n=n^{\prime}+p$, where $p$ represents the additional parameters). In the case of such nested models, under certain mild assumptions, $\ln B$ can be analytically expressed and depends not only on the difference in chi-squares of the two models but also on the Fisher matrix of the additional parameters and their priors. In our case, the latter two factors may be what contribute to the observed difference.
We have also performed a comparison with a two parameter extension of the $\Lambda$CDM model, the so-called $w_{0}w_{a}$, and summarize the results in Appendix B.